# Subfitness in distributive (semi)lattices

### v1

@@ 1. Introduction
{tikzpicture}
At the BLAST conference at Chapman University in August 2022, coauthor Madden asked: Do the join-subfit elements of a distributive join-semilattice form an ideal? It is easy to show that in a join-semilattice, if $a\leq b$ and $b$ is join-subfit, then so is $a$. So the problem is to show that if $a$ and $b$ are join-subfit, then so is $a\vee b$; in other words, to show that the join-subfit elements of $A$ are directed. A simple example shows that the distributive hypothesis is necessary; in the lattice below, $a$ and $b$ are join-subfit, but $a\vee b$ is not.
{tikzpicture}
At the time it was asked, the answer was known to be “Yes” for finite semilattices. The question is motivated by the following considerations. If a topological space is the union of two subspaces, both open and regular, the space may fail to be regular – the line with a point doubled is the classical example, see [Mun00, Chap. 4, Ex. 5, p. 227]. It is meaningful to ask if similar examples arise if we take subfitness in place of regularity. The question arises in several more general contexts related to representations of algebras (see [Joh62]). A related question was asked in [DIM21]: if $A$ is a join-semilattice such that ${\downarrow}a$ is join-subfit for each $a\in A$, is $A$ ideally subfit? [^2] The property of ideal subfitness was introduced in [MZ06] and studied in [DIM21]. A join-semilattice $S$ is said to be ideally subfit if for all $u,v\in A$, if $u\nleq v$, then there is an ideal $W$ of $A$ such that, in the lattice of ideals of $A$, we have ${\downarrow}u\vee W=A$ and ${\downarrow}v\vee W\neq A$. Considering the relationship to the regularity problem, it is reasonable to ask whether the directedness condition is necessary, and this is the immediate motivation for the question.
@@ 3. Subfit elements in distributive semilattices are not directed
An illustration is given in Fig. 1, where the cup shaped bubbles represent sets of finite sets and the unique umbrella shape at the top is a set of cofinite sets. We set $a=\{0\},\,b=\{1\},\,c=\{1,2\}$, and $\mathcal{F}_{012}:=\{\,F\cup\{0,1,2\}\mid F\in\mathcal{F}\,\}$. Then $a,b,c\in A$ and $A$ is depicted in the black part of the figure. The set $\mathcal{F}_{012}$ is disjoint from $A$, but is contained in the sublattice of $\mathcal{P}(\mathbb{N})$ generated by $A$.
{tikzpicture}
@@
Suppose $x,y\in{\uparrow}a$ and $y\nsubseteq x$. We need to find $z\in{\uparrow}a$ with $x\cap z=\varnothing$ and $y\cap z\neq\varnothing$. Pick $n\in y\setminus x$. (Of course, $n\neq 0$.) If $n\neq 2$, let $z=\{0,n\}$. If $n=2$, then $y$ is cofinite and $x$ is finite, unless it contains 2. Therefore, $y\setminus x$ contains an integer $m\geq 3$, and we set $z=\{0,m\}$. ∎
@@
Consider $y=\{1,2\}$ (i.e. $y=c$) and let $x$ be any element of $A$ that contains $1$ but not $2$. The only non-zero elements of $A$ that are less than or equal to $y$ are $y$, $\{1\}$ and $\varnothing$, so the desired $z$ does not exist. ∎
@@
    Suppose $z\in\mathcal{F}_{01}$. In this case $u\cup z\in A$ and we argue as in (i), unless $u\in\mathcal{F}_{12}$. If $x\in\mathcal{F}_{12}$, then $y\in\mathcal{F}\cup\mathcal{F}_{0}\cup\mathcal{F}_{01}$ and we proceed as in (ii).
@@ 4. Distributive envelopes and subfitness
In this section we show that to each join semilattice $A$ – distributive or not – one can associate a distributive lattice $L$ such that $A$ is join-subfit iff $L$ is join-subfit (and analogously for meets); see Theorem 4.2. The construction of $L$ from $A$ is functorial and will be studied in a forthcoming paper. If $A$ happens to be weakly distributive in the sense of [Bal69, Var75], then $L$ is the distributive envelope of $A$ defined in [CH78] (and also studied in [HP08, BJ11]); we won’t use this explicitly but leave comments en route to show the connection. Instead we will take $L$ as the sublattice generated by $A$ in its injective hull following [BL70, HK71] in its incarnation for join-semilattices.
@@
    $E$ is a distributive lattice.555in fact it is a co-frame, i.e. the order-dual of a frame.
@@
Let $A$ be a bounded join-semilattice and $E$ be the injective hull of $A$. We set $L$ to be the bounded sublattice of $E$ generated by $A$. Then $L$ is a bounded distributive lattice and $A$ is join-subfit if and only if $L$ is join-subfit.∎
@@ 5. Topological perspective
[…]
@@
    $X$ is compact if and only if ${\mathrm{cp}(X)}$ is compact and every nonempty closed subset of $X$ contains a closed point, cf. [DST19, 4.1.2]
@@
First suppose ${\mathrm{cp}(X)}$ is dense for the patch-topology of $X$. Take $U,V\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $U\nsubseteq V$ and take a closed point $x\in U\setminus V$. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis and $p$ is a closed point, $(X\setminus U)\cup V\subseteq X\setminus\{p\}=\bigcup\{W:p\notin W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)\}.$ Because $(X\setminus U)\cup V$ is compact and $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is closed under finite unions, there is $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $(X\setminus U)\cup V\subseteq W\not\ni p$. Thus, $p\notin W\cup V$ and $U\cup W=X$, showing that $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit.
Conversely assume that $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis, the sets of the form $U\setminus V$ with $U,V\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ form a basis for the patch-topology of $X$. So we only need to show that each nonempty such set has a closed point. As $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit, there is $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $U\cup W=X$ and $V\cup W\neq X$. This means $\varnothing\neq X\setminus W\subseteq U\setminus W$. Now $X\setminus W$ is a nonempty closed set of the compact space $X$, hence it contains a closed point of $X$ by Remark 5.1ii. ∎
@@
  1. (*)(*)( * )
    If $O_{1},O_{2}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $O_{1}\subseteq O_{2}$, then for every patch open subset $C$ of the space $O_{2}$, the set $C\cap O_{1}$ is empty or has nonempty interior for the patch topology of $O_{1}$. [^7] In fact, we even know that the patch topology of $O_{1}$ is the subspace topology induced on $O_{1}$ by the patch topology of $O_{2}$. The weaker assumption (*)(*)( * ) is used here because the proof also goes through for compact, compactly based spaces only satisfying (*)(*)( * ).
We need to show that ${\mathrm{cp}(U\cup V)}$ is patch-dense in $U\cup V$. Take $C\subseteq U\cup V$ nonempty and patch-open. By symmetry we may assume that $C\cap V\neq\varnothing$. By (*)(*)( * ) and the assumption on $V$, there is some $v\in{\mathrm{cp}(V)}\cap C$. If $v\notin U$, then since $U$ is open, $v$ is a closed point of $U\cup V$ and we are done. Hence, we may assume that $v\in U$.
From $v\in{\mathrm{cp}(V)}$ we know $\overline{\{v\}}\cap V=\{v\}$ and so ${\mathrm{cp}(V)}\setminus U\subseteq X\setminus\overline{\{v\}}$. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis and ${\mathrm{cp}(V)}\setminus U$ is compact (use Remark 5.1ii), there is some $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with ${\mathrm{cp}(V)}\setminus U\subseteq W\subseteq X\setminus\overline{\{v\}}$. Consequently, $v\in U\cap C\setminus W$, showing that $U\cap C\setminus W$ is nonempty. By (*)(*)( * ) and the assumption on $U$, there is some $u\in{\mathrm{cp}(U)}\cap C\setminus W$. If $u\notin V$, then since $V$ is open, $u\in{\mathrm{cp}(U\cup V)}$ and we are done. Hence, we may assume that $u\in V$. If $u\in{\mathrm{cp}(V)}$, then we are done. Otherwise there is $w\in{\mathrm{cp}(V)}$ such that $w\in\overline{\{u\}}$ and $w\neq u$. But then $w\notin U$ because $u\in{\mathrm{cp}(U)}$. Thus, $w\in{\mathrm{cp}(V)}\setminus U\subseteq W$ and so $u\in W$, a contradiction. ∎
@@
{tikzpicture}
Here $\omega$ is the only non-isolated point of the patch topology of $Y$ and the immediate specializations are those indicated by a squiggly arrow [^8] Hence, $\omega\rightsquigarrow x$ means that $\overline{\{\omega\}}$ is the only closed subset of $\overline{\{\omega\}}$ containing $x$. (hence all points of $Y$ except $\omega$ and $z$ are closed points). Let $X$ be the subspace $Y\setminus\{\omega\}$ of $Y$ and let $P=\{p_{0},p_{1},p_{2},\ldots\}$. One checks that for $U\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(Y)$ the set $U\cap X$ is compact provided $U\neq P_{0}\cup\{\omega\}$ for all cofinite subsets $P_{0}$ of $P$. It follows that $X$ is compactly based. However $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is not join-subfit by Proposition 5.2 since ${\mathrm{cp}(X)}$ is not patch-dense, witnessed by the identity $\{z\}=(P\cup\{z\})\setminus(P\cup\{x\})$.
Finally, $X=V\cup W$ with $V=P\cup\{x\}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ and $W=P\cup\{y,z\}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$, but $\mathop{\mathring{{\mathcal{K}}}}\nolimits(V)$ and $\mathop{\mathring{{\mathcal{K}}}}\nolimits(W)$ are join-subfit: One checks that $V,W$ are even spectral and their subsets of closed points are patch-dense.

### v2

@@ 1. Introduction
A topology, seen as a join-semilattice, is join-subfit iff every nonempty locally closed set contains a nonempty closed set (cf. [PP21, p. 23]). This shows that join-subfitness of a topology is near the T1 separation property. In fact, a topology is $T_{1}$ iff it is $T_{D}$ and join-subfit; see [PP21, p. 24]. Subfitness also has a prominent role in algebraic geometry: If $X$ is the prime spectrum of a commutative ring $R$, then its frame of radical ideals is join-subfit if and only if $R$ is a Jacobson ring, i.e. every radical ideal of $R$ is the intersection of maximal ideals of $R$. Prominent examples are finitely generated algebras over a given Jacobson ring; every homomorphism $\phi:R\longrightarrow S$ of such algebras induces a map between the maximal ideal spectra $\mathrm{Max}(S)\longrightarrow\mathrm{Max}(R),{\mathfrak{m}}\mapsto\phi^{-1}({\mathfrak{m}})$ (see [DST19, Thm. 12.3.12 and Rem. 12.3.14]). The significance of join-subfitness for ring theory in connection with Jacobson rings was pointed out by Simmons [Sim78] where the property was named conjunctive; see also [DIM21].
$a$$b$$a^{\prime}$$c$$b^{\prime}$$a\vee b$
At the BLAST conference at Chapman University in August 2022, coauthor Madden asked: Do the join-subfit elements of a distributive join-semilattice form an ideal? It is easy to show that in a distributive join-semilattice, if $a\leq b$ and $b$ is join-subfit, then so is $a$. So the problem is to show that if $a$ and $b$ are join-subfit, then so is $a\vee b$; in other words, to show that the join-subfit elements of $A$ are directed. A simple example shows that the distributive hypothesis is necessary; in the lattice below, $a$ and $b$ are join-subfit, but $a\vee b$ is not.
@@
At the time it was asked, the answer was known to be “Yes” for finite semilattices. The question is motivated by the following considerations. If a topological space is the union of two subspaces, both open and regular, the space may fail to be regular – the line with a point doubled is the classical example, see [Mun00, Chap. 4, Ex. 5, p. 227]. It is meaningful to ask if similar examples arise if we take subfitness in place of regularity. The question arises in several more general contexts related to representations of algebras (see [Joh62]). A related question was asked in [DIM21]: if $A$ is a join-semilattice such that ${\downarrow}a$ is join-subfit for each $a\in A$, is $A$ ideally subfit? [^2] The property of ideal subfitness was introduced in [MZ06] and studied in [DIM21]. A join-semilattice $S$ is said to be ideally subfit provided for all $u,v\in A$, if $u\nleq v$, then there is an ideal $W$ of $A$ such that, in the lattice of ideals of $A$, we have ${\downarrow}u\vee W=A$ and ${\downarrow}v\vee W\neq A$. Considering the relationship to the regularity problem, it is reasonable to ask whether the directedness condition is necessary, and this is the immediate motivation for the question.
@@ 3. Subfit elements in distributive semilattices are not directed
An illustration is given in Fig. 1, where the cup shaped bubbles represent sets of finite sets and the unique umbrella shape at the top is a set of cofinite sets. We set $a:=\{0\},\,b:=\{1\},\,c:=\{1,2\}$, and $\mathcal{F}_{012}:=\{\,F\cup\{0,1,2\}\mid F\in\mathcal{F}\,\}$. Then $a,b,c\in A$ and $A$ is depicted in the black part of the figure. The set $\mathcal{F}_{012}$ is disjoint from $A$, but is contained in the sublattice of $\mathcal{P}(\mathbb{N})$ generated by $A$.
$\scriptstyle\varnothing\scriptstyle\mathcal{F}$$\scriptstyle a$$\scriptstyle\mathcal{F}_{0}$$\scriptstyle b$$\scriptstyle\mathcal{F}_{1}$$\scriptstyle a\vee b$$\scriptstyle\mathcal{F}_{01}$$\scriptstyle c$$\scriptstyle\mathcal{F}_{12}$$\scriptstyle a\vee c$$\scriptstyle\mathcal{F}_{012}$$\scriptstyle\mathcal{C}_{012}$$\scriptstyle\mathbb{N}$
@@
Suppose $x,y\in{\uparrow}a$ and $y\nsubseteq x$. We need to find $z\in{\uparrow}a$ with $x\cap z=a$ and $y\cap z\neq a$. Pick $n\in y\setminus x$. (Of course, $n\neq 0$.) If $n\neq 2$, let $z=\{0,n\}$. If $n=2$, then $y$ is cofinite and $x$ is finite. Therefore, $y\setminus x$ contains an integer $m\geq 3$, and we set $z=\{0,m\}$. ∎
@@
Consider $y=\{1,2\}$ (i.e. $y=c$) and let $x$ be any element of $A$ that contains $1$ but not $2$. Then $y\nsubseteq x$, but the only non-zero elements of $A$ that are less than or equal to $y$ are $y$, $\{1\}$ and $\varnothing$, so the desired $z$ does not exist. ∎
@@
    Suppose $z\in\mathcal{F}_{01}$. In this case $u\cup z\in A$ and we argue as in (i), unless $u\in\mathcal{F}_{12}$. If $x\in\mathcal{F}_{12}$, then $y\in\mathcal{F}\cup\mathcal{F}_{0}\cup\mathcal{F}_{1}\cup\mathcal{F}_{01}$ and we proceed as in (ii).
@@ 4. Distributive envelopes and subfitness
In this section we show that to each join-semilattice $A$ – distributive or not – one can associate a distributive lattice $L$ such that $A$ is join-subfit iff $L$ is join-subfit (and analogously for meets); see Theorem 4.2. The construction of $L$ from $A$ is functorial and will be studied in a forthcoming paper. If $A$ happens to be weakly distributive in the sense of [Bal69, Var75], then $L$ is the distributive envelope of $A$ defined in [CH78] (and also studied in [HP08, BJ11]); we won’t use this explicitly but leave comments en route to show the connection. Instead we will take $L$ as the sublattice generated by $A$ in its injective hull following [BL70, HK71] in its incarnation for join-semilattices.
@@
    $E$ is a distributive lattice. [^5] In fact it is a co-frame, i.e. the order-dual of a frame.
@@
Let $A$ be a bounded join-semilattice and let $E$ be the injective hull of $A$. We set $L$ to be the bounded sublattice of $E$ generated by $A$. Then $L$ is a bounded distributive lattice and $A$ is join-subfit if and only if $L$ is join-subfit.∎
@@ 5. Topological perspective
Join-semilattices occur naturally in non-Hausdorff topology. For a topological space $X$ (which always means T0 here), the set
@@
    $X$ is compact if and only if ${\mathrm{cp}(X)}$ is compact and every nonempty closed subset of $X$ contains a closed point, cf. [DST19, 4.1.2].
@@
First suppose ${\mathrm{cp}(X)}$ is dense for the patch-topology of $X$. Take $U,V\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $U\nsubseteq V$ and take a closed point $p\in U\setminus V$. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis and $p$ is a closed point, $(X\setminus U)\cup V\subseteq X\setminus\{p\}=\bigcup\{W:p\notin W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)\}.$ Because $(X\setminus U)\cup V$ is compact and $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is closed under finite unions, there is $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $(X\setminus U)\cup V\subseteq W\not\ni p$. Thus, $p\notin W\cup V$ and $U\cup W=X$, showing that $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit.
Conversely, assume that $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis, the sets of the form $U\setminus V$ with $U,V\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ form a basis for the patch-topology of $X$. So we only need to show that each nonempty such set has a closed point. As $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is join-subfit, there is $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $U\cup W=X$ and $V\cup W\neq X$. This implies $\varnothing\neq X\setminus(V\cup W)\subseteq U\setminus V$. Now, $X\setminus(V\cup W)$ is a nonempty closed set of the compact space $X$, hence it contains a closed point of $X$ by Remark 5.1ii. ∎
@@
  1. ()(*)( ∗ )
    If $O_{1},O_{2}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with $O_{1}\subseteq O_{2}$, then for every patch open subset $C$ of the space $O_{2}$, the set $C\cap O_{1}$ is empty or has nonempty interior for the patch topology of $O_{1}$. [^7] In fact, we even know that the patch topology of $O_{1}$ is the subspace topology induced on $O_{1}$ by the patch topology of $O_{2}$. The weaker assumption ()(*)( ∗ ) is used here because the proof also goes through for compact, compactly based spaces only satisfying ()(*)( ∗ ).
We need to show that ${\mathrm{cp}(U\cup V)}$ is patch-dense in $U\cup V$. Take $C\subseteq U\cup V$ nonempty and patch-open. By symmetry we may assume that $C\cap V\neq\varnothing$. By ()(*)( ∗ ) and the assumption on $V$, there is some $v\in{\mathrm{cp}(V)}\cap C$. If $v\notin U$, then since $U$ is open, $v$ is a closed point of $U\cup V$ and we are done. Hence, we may assume that $v\in U$.
From $v\in{\mathrm{cp}(V)}$ we know $\overline{\{v\}}\cap V=\{v\}$ and so ${\mathrm{cp}(V)}\setminus U\subseteq X\setminus\overline{\{v\}}$. Since $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is a basis and ${\mathrm{cp}(V)}\setminus U$ is compact (use Remark 5.1ii), there is some $W\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ with ${\mathrm{cp}(V)}\setminus U\subseteq W\subseteq X\setminus\overline{\{v\}}$. Consequently, $v\in U\cap C\setminus W$, showing that $U\cap C\setminus W$ is nonempty. By ()(*)( ∗ ) and the assumption on $U$, there is some $u\in{\mathrm{cp}(U)}\cap C\setminus W$. If $u\notin V$, then since $V$ is open, $u\in{\mathrm{cp}(U\cup V)}$ and we are done. Hence, we may assume that $u\in V$. If $u\in{\mathrm{cp}(V)}$, then we are done. Otherwise there is $w\in{\mathrm{cp}(V)}$ such that $w\in\overline{\{u\}}$ and $w\neq u$. But then $w\notin U$ because $u\in{\mathrm{cp}(U)}$. Thus, $w\in{\mathrm{cp}(V)}\setminus U\subseteq W$ and so $u\in W$, a contradiction. ∎
@@
$p_{0}$$p_{1}$$p_{2}$$p_{3}$$\omega$$x$$z$$y$
Here $\omega$ is the only non-isolated point of the patch topology of $Y$ and the immediate specializations are those indicated by a squiggly arrow [^8] Hence $u\rightsquigarrow v$ implies that $v\in\overline{\{u\}}$. (hence all points of $Y$ except $\omega$ and $z$ are closed points). Let $X$ be the subspace $Y\setminus\{\omega\}$ of $Y$ and let $P=\{p_{0},p_{1},p_{2},\ldots\}$. One checks that for $U\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(Y)$ the set $U\cap X$ is compact provided $U\neq P_{0}\cup\{\omega\}$ for all cofinite subsets $P_{0}$ of $P$. It follows that $X$ is compactly based. However $\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ is not join-subfit by Proposition 5.2 since ${\mathrm{cp}(X)}$ is not patch-dense, witnessed by the identity $\{z\}=(P\cup\{z\})\setminus(P\cup\{x\})$.
Finally, $X=V\cup W$ with $V=P\cup\{x\}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$ and $W=P\cup\{y,z\}\in\mathop{\mathring{{\mathcal{K}}}}\nolimits(X)$, but $\mathop{\mathring{{\mathcal{K}}}}\nolimits(V)$ and $\mathop{\mathring{{\mathcal{K}}}}\nolimits(W)$ are join-subfit as one checks using Proposition 5.2 and the characterization of compactness above.